# Exploring and Leveraging Class Vectors for Classifier Editing

### v1

@@ C.2 Existence of a Mapping Function
For any input $x$, if the editable-parameter perturbation $w$ is sufficiently small, then
@@
so that a first-order Taylor approximation around $\theta$ is valid.
Let $J:=J_{f}(\theta)\in\mathbb{R}^{m\times d_{e}}$ denote the Jacobian of the encoder output with respect to the editable parameters, evaluated at the current weights $\theta$. Because edits are restricted to an overparameterised subset of layers with $d_{e}\gg m$, we may assume that $\operatorname{rank}(J)=m$ (full row rank).
Step 1 (solving a linear constraint).
For any target vector $z\in\mathbb{R}^{m}$ we require a perturbation $w\in\mathbb{R}^{d_{e}}$ with
Jw=z.J\,w\;=\;z.
Since $J$ has full row rank, the matrix $JJ^{\top}\in\mathbb{R}^{m\times m}$ is invertible and the Moore–Penrose pseudoinverse $J^{\dagger}:=J^{\top}(JJ^{\top})^{-1}\in\mathbb{R}^{d_{e}\times m}$ satisfies $JJ^{\dagger}=I_{m}$. Hence $w_{0}(z):=J^{\dagger}z$ is one solution of (A).
Step 2 (characterising all solutions).
Define the null space $\mathcal{N}:=\ker J=\{v\in\mathbb{R}^{d_{e}}\mid Jv=0\}.$ Its dimension is $d_{e}-m>0$, so $\mathcal{N}$ contains non-zero vectors. For any $v\in\mathcal{N}$ the vector $w(z):=w_{0}(z)+v$ also satisfies (A). Therefore the full solution set is the affine subspace
$\mathcal{S}(z)\;=\;J^{\dagger}z\;+\;\ker J,$
which has dimension $d_{e}-m\geq 1$. Consequently, there are infinitely many distinct solutions for every $z$.
Step 3 (building infinitely many mappings).
Choose any linear map $N\in\mathbb{R}^{d_{e}\times m}$ whose columns lie in $\ker J$ ( $JN=0$ ). Then the matrix
$R\;:=\;J^{\dagger}+N$
is a right inverse of $J$ because $JR=JJ^{\dagger}=I_{m}$. Define
$\phi_{\mathrm{edit}}(z)\;:=\;Rz\qquad(z\in\mathbb{R}^{m}).$
By construction $J\,\phi_{\mathrm{edit}}(z)=z$ for all $z$. Since $\ker J$ is non-trivial, we can vary $N$ (or add any vector in $\ker J$ to $Rz$) to obtain infinitely many distinct mappings that all satisfy the required constraint.
Step 4 (smallness of the perturbation).
Because $\ker J$ is a linear subspace, for any $\varepsilon>0$ we can scale the target $z$ (or equivalently scale $R$) so that $\|\phi_{\mathrm{edit}}(z)\|\leq\varepsilon$, ensuring the weight perturbation remains sufficiently small while still realising the desired latent edit.
Hence the conditions of the theorem guarantee the existence of infinitely many weight-space mappings $\phi_{\mathrm{edit}}:\mathbb{R}^{m}\to\mathbb{R}^{d_{e}}$. ∎

### v2

@@ C.2 Existence of a Mapping Function
For any input $x$, if the editable–parameter perturbation $w$ is sufficiently small, then
@@
so a first–order Taylor approximation around $\theta$ is valid. Let $J:=J_{f}(\theta)\in\mathbb{R}^{m\times d_{e}}$ denote the Jacobian of the encoder output with respect to the editable parameters, evaluated at $\theta$. Since edits are restricted to an overparameterised subset of layers with $d_{e}\gg m$, we assume $\operatorname{rank}(J)=m$ (full row rank).
We seek a perturbation $w\in\mathbb{R}^{d_{e}}$ that realises a target change $z\in\mathbb{R}^{m}$ in the encoder output, i.e.
$Jw=z.$
Because $J$ has full row rank, the matrix $JJ^{\top}\in\mathbb{R}^{m\times m}$ is invertible, and the Moore–Penrose pseudoinverse
$J^{\dagger}:=J^{\top}(JJ^{\top})^{-1}\in\mathbb{R}^{d_{e}\times m}$
satisfies $JJ^{\dagger}=I_{m}$. Hence one particular solution of $(A)$ is
$w_{0}(z):=J^{\dagger}z.$
Let $\mathcal{N}:=\ker J=\{v\in\mathbb{R}^{d_{e}}\mid Jv=0\}$. Its dimension is $d_{e}-m>0$, so the full solution set of $(A)$ is the affine subspace
$\mathcal{S}(z)=J^{\dagger}z+\ker J=\{\,J^{\dagger}z+v\mid v\in\ker J\,\}.$
Since $\mathcal{N}$ is nontrivial, infinitely many distinct $w$ satisfy $Jw=z$.
Now take any linear map $N\in\mathbb{R}^{d_{e}\times m}$ whose columns lie in $\ker J$ ($JN=0$), and define
$R:=J^{\dagger}+N.$
Then $JR=JJ^{\dagger}=I_{m}$, so for every $z\in\mathbb{R}^{m}$,
$J(Rz)=z.$
The mapping
$\phi_{\mathrm{edit}}(z):=Rz$
thus provides a valid weight–space perturbation realising the desired edit. Varying $N$ (or equivalently adding any vector in $\ker J$ to $Rz$) yields infinitely many distinct mappings $\phi_{\mathrm{edit}}:\mathbb{R}^{m}\to\mathbb{R}^{d_{e}}$ that all satisfy $J\,\phi_{\mathrm{edit}}(z)=z$.
Finally, since $\ker J$ is a linear subspace, for any $\varepsilon>0$ we may rescale the target $z$ (or equivalently $R$) to ensure $\|\phi_{\mathrm{edit}}(z)\|\leq\varepsilon$, so the perturbation remains sufficiently small while achieving the desired first–order effect. Therefore, under the stated conditions, there exist infinitely many sufficiently small weight–space mappings $\phi_{\mathrm{edit}}$ that satisfy $J\,\phi_{\mathrm{edit}}(z)=z$ for all $z\in\mathbb{R}^{m}$. ∎